You need a new cell phone for emergencies only company a charges $12 per month plus $0.10 per minute while Company B charges $0.15 per minute with no monthly service charge for how many minutes will the monthly cost to be the same
Found 2 solutions by rfer, addingup:
Answer by rfer(16322)

12+0.10x=0.15x
12=0.15x-0.10x
12=0.05x
12/0.05=x
x=240

Answer by addingup(3677)

Let's call the minutes x:
Company A:
12+0.10x
Company B:
0.15x When will costs be the same? When A=B:
12+0.10x=0.15x Now subtract 0.10x from both sides:
12= 0.05x
Divide both sides by 0.05:
x=240 At 240 minutes both will cost you the same.
--------------------------
Now, every minute OVER 240 will cost you 0.10 with A and 0.15 with B, therefore if you are going to use more than 240 minutes choose A.
On the other hand, if you are going to use LESS than 240 minutes... what happens? Let's pretend you will use 200 minutes:
A)12+0.10(200)= 32
B)0.15(200)= 30
So, as you can see, if you use less than 240 minutes B is cheaper. k ijeoi